What is the density of a sample of hydrogen gas (H2) that exerts 2.85 atm of pressure at 225K?

  • Use 0.08206L atmmol K for the ideal gas constant.
  • Round the answer to three significant figures.
Answer #1

P = 2.85atm

T = 225 K

Molar mass of H2 = 2.016 g/mol

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

2.85 atm *2.016 g/mol = density * 0.08206 atm.L/mol.K *225.0 K

density = 0.3112 g/L

Answer: 0.311 g/L